# Privacy of the last iterate in cyclically-sampled DP-SGD on nonconvex composite losses

### v1

@@ Appendix A Technical proofs
We are now ready to give the main proof of this subsection.
Proof of Lemma 2.1..
@@
$\leq D_{\alpha}^{(L\tau+s)}(\phi(Y)\|\phi^{\prime}(Y^{\prime}))+\frac{\alpha a^{2}}{2\sigma^{2}}\leq D_{\alpha}^{(\tau)}(Y\|Y^{\prime})+\frac{\alpha a^{2}}{2\sigma^{2}}.\qed$
A.2 Proof of Proposition 2.2
We start with a technical lemma.
@@
$\sum_{t=1}^{T}b_{t}^{2}$ $=\frac{L^{2(T-1)}\sum_{i=0}^{T-1}L^{2i}}{\left(\sum_{i=0}^{T-1}L^{2i}\right)^{2}}=\frac{L^{2(T-1)}}{S_{T}}=\theta_{L}(T).\qed$
The next result considers the case of a single dataset pass with $T=\ell$.
@@
Let $L$, $\tau$, $\zeta$, $\bar{s}$, $\{Y_{t}\}$, $\{Y_{t}^{\prime}\}$, $\ell$, and $T$ be as in Proposition 2.2. If $T=\ell$, then the following statements hold:
@@
Using Lemma A.2(a)–(b), we first have $L\tau+s-a_{1}=(L\tau+s)R_{1}\geq 0$ and, hence, by Lemma 2.1, we have
@@
Since Lemma A.2(a)–(b) also implies $R_{t}\geq 0$ and we have $s_{t}=0$ for $t\geq 2$, we repeatedly apply Lemma 2.1 with $(a,\tau)=(a_{t},\tau_{t})=(a_{t},0)$ for $t\geq 2$ to obtain
@@
(b) Let $s=\bar{s}$. Similar to (a), our goal is to recursively apply (5) with suitable choices of the free parameter $a$ at each application. For this setting, let $a_{1}=\zeta+s$ and $a_{t}=\zeta$ for $t\geq 2$. Using the fact that $\tau=0$ and $L=1$ and Lemma 2.1, we first have that
@@
We then repeatedly apply Lemma 2.1 with $(a,\tau)=(a_{t},0)$ for $t\geq 2$ to obtain
@@
We are now ready to give the main proof of this subsection.
Proof of Proposition 2.2.
@@
${\cal B}_{1}(\tau,T)$ $:=\frac{\alpha}{2}\left(\frac{L\tau+s}{\sigma}\right)^{2}\theta_{L}(T),$
${\cal B}_{2}$ $:=\frac{\alpha}{\sigma^{2}}\left[(L\tau+s)^{2}+s^{2}\left\{(E-1)\theta_{L}(\ell)+\theta_{L}(T-E\ell)\right\}\right].$
@@
$D_{\alpha}(Y_{T}\|Y_{T}^{\prime})-D_{\alpha}(Y_{0}\|Y_{0}^{\prime})$
$\leq{\cal B}_{1}(\tau,\ell)+(E-1){\cal B}_{1}(0,\ell)+{\cal B}_{1}(0,T-E\ell)={\cal B}_{2}.$
@@
$D_{\alpha}(Y_{T}\|Y_{T}^{\prime})$ $\leq D_{\alpha}(Y_{E\ell}\|Y_{E\ell}^{\prime})+2\alpha[T-E\ell]\left(\frac{\zeta}{\sigma}\right)^{2}\leq D_{\alpha}(Y_{0}\|Y_{0}^{\prime})+2\alpha T\left(\frac{\zeta}{\sigma}\right)^{2}.\qed$
A.3 Proof of Theorem 4.5
The below result gives a general divergence bound that primarily depends on $i^{*}$, $E$, and $\ell$.
Proposition A.4.
Let $\{X_{t}\}$, $\{X_{t}^{\prime}\}$, $\lambda$, $b$, $C$, and $T$ be as in Theorem 4.4, and let $t^{*}:=\inf_{t\geq 1}\left\{t:i^{*}\in B_{t}\right\}$. If $\lambda\leq 1/[2(m+M)]$ and $T\geq t^{*}+1$, then
$D_{\alpha}(X_{T}\|X_{T}^{\prime})$ $\leq 2\alpha\left(\frac{\lambda C}{b\sigma}\right)^{2}\left[E\theta_{L_{\lambda}}(\ell)+\theta_{L_{\lambda}}(T-t^{*}-1-E\ell)\right],$ (27)
where $E:=\lfloor(T-t^{*}-1)/\ell\rfloor$, $\ell:=k/b$, and $L_{\lambda}$ is as in (12).
@@
Using Lemma 4.2(b) and Proposition 2.2(a) with
@@
$=2\alpha\left(\frac{\lambda C}{b\sigma}\right)^{2}\left[E\theta_{L_{\lambda}}(\ell)+\theta_{L_{\lambda}}(T-t^{*}-1-E\ell)\right].\qed$
We are now ready to give the main proof of this subsection.
Proof of Theorem 4.5.
The bound in (21) follows from Proposition A.4, the fact that $t^{*}\geq 1$, and the fact that $\theta_{L_{\lambda}}(s)\leq 1$ for every $s\geq 1$, and the fact that $\theta_{L_{\lambda}}(s)$ is monontonically decreasing for $s\geq 1$. Now, assume that each $f_{i}$ is convex. The fact that $m=0$ is well-known. Next, observe that the definition of $L_{\lambda}=L_{\lambda}(m,M)$ in (12) implies that $L_{\lambda}(0,M)=1$ and the definition of $\theta_{L}(\cdot)$ in (6) implies $\theta_{1}(s)=1/s$ for every $s\geq 1$ and $\theta_{1}(0)=0$. The bound in (22) now follows from using the previous identities for $\theta_{1}(\cdot)$ in (21). ∎

### v3

@@ A.1 General operator analysis
The next result is the aforemention shifted RDP bound.
Lemma A.2.
For some $L,\zeta\geq 0$, suppose $\phi^{\prime}$ and $\phi$ satisfy (8) for any $x,y\in\operatorname*{dom}\phi\cap\operatorname*{dom}\phi^{\prime}$. Moreover, let $Z\sim{\cal N}(0,\sigma^{2}I)$ and $\psi\in\overline{{\rm Conv}}\,(\mathbb{R}^{n})$. Then, for any scalars $a,\tau\geq 0$ and $\alpha\in(1,\infty)$ satisfying $L\tau+\zeta+s-a\geq 0,$ and random variables $Y$ and $Y^{\prime}$, it holds that
Proof.
@@
$\leq D_{\alpha}^{(L\tau+s)}(\phi(Y)\|\phi^{\prime}(Y^{\prime}))+\frac{\alpha a^{2}}{2\sigma^{2}}\leq D_{\alpha}^{(\tau)}(Y\|Y^{\prime})+\frac{\alpha a^{2}}{2\sigma^{2}}.$
Note that the second inequality in (8) is equivalent to $\Phi$ being $L$-Lipschitz continuous when $\zeta=0$, and that the conditions in (8) need to only hold on $\operatorname*{dom}\phi\cap\operatorname*{dom}\phi^{\prime}$.
We next apply (20) to a sequence of points generated by the update
$Y\xleftarrow{}\operatorname{prox}_{\psi}(\phi(Y)+Z)$ (22)
under different assumptions on $\zeta$ and $\tau$ and a single dataset pass. Before proceeding, we first present a technical lemma.
@@
$\sum_{t=1}^{T}b_{t}^{2}$ $=\frac{L^{2(T-1)}\sum_{i=0}^{T-1}L^{2i}}{\left(\sum_{i=0}^{T-1}L^{2i}\right)^{2}}=\frac{L^{2(T-1)}}{S_{T}}=\theta_{L}(T).$
We now present the shifted RDP properties of the update in (22). This particular result generalizes the one in [17], which only considers the case of $L=1$ and $\zeta=0$.
@@
Let $L,\zeta\geq 0$, $T\geq 1$, and $\ell\in[T]$ be fixed. Given $\psi\in\overline{{\rm Conv}}\,(\mathbb{R}^{n})$, suppose $\{\phi_{t}\}_{t=1}^{T}$, $\{\phi_{t}^{\prime}\}_{t=1}^{T}$, and $\bar{s}>0$ satisfy (8) with
$\begin{gathered}Y_{t}:=\operatorname{prox}_{\psi}(\phi_{t}(Y_{t-1})+Z_{t}),\quad Y_{t}^{\prime}:=\operatorname{prox}_{\psi}(\phi_{t}^{\prime}(Y_{t-1}^{\prime})+Z_{t}),\quad\forall t\geq 1.\end{gathered}$
If $T=\ell$, then the following statements hold:
@@
Using Lemma A.3(a)–(b), we first have $L\tau+s-a_{1}=(L\tau+s)R_{1}\geq 0$ and, hence, by Lemma A.2, we have
@@
Since Lemma A.3(a)–(b) also implies $R_{t}\geq 0$ and we have $s_{t}=0$ for $t\geq 2$, we repeatedly apply Lemma A.2 with $(a,\tau)=(a_{t},\tau_{t})=(a_{t},0)$ for $t\geq 2$ to obtain
@@
(b) Let $s=\bar{s}$. Similar to (a), our goal is to recursively apply (20) with suitable choices of the free parameter $a$ at each application. For this setting, let $a_{1}=\zeta+s$ and $a_{t}=\zeta$ for $t\geq 2$. Using the fact that $\tau=0$ and $L=1$ and Lemma A.2, we first have that
@@
We then repeatedly apply Lemma A.2 with $(a,\tau)=(a_{t},0)$ for $t\geq 2$ to obtain
@@
We next extend the above result to multiple dataset passes.
Proposition A.5.
Let $L$, $\tau$, $\zeta$, $\bar{s}$, $\{Y_{t}\}$, $\{Y_{t}^{\prime}\}$, $\ell$, and $T$ be as in Lemma A.4. Moreover, let $\theta_{L}(\cdot)$ be as in (15). For any $\tau\geq 0$ and $E=\lfloor T/\ell\rfloor$, the following statements hold:
  • (a)
    if $\zeta=0$, then
    $D_{\alpha}(Y_{T}\|Y_{T}^{\prime})-D_{\alpha}^{(\tau)}(Y_{0}\|Y_{0}^{\prime})$
  • (b)
    if $\tau=0$ and $\zeta=\bar{s}$, then
Proof.
@@
${\cal B}_{1}(\tau,T)$ $:=\frac{\alpha}{2}\left(\frac{L\tau+s}{\sigma}\right)^{2}\theta_{L}(T),\quad{\cal B}_{2}:=\frac{\alpha}{\sigma^{2}}\left[(L\tau+s)^{2}+s^{2}\left\{(E-1)\theta_{L}(\ell)+\theta_{L}(T-E\ell)\right\}\right].$
@@
$D_{\alpha}(Y_{T}\|Y_{T}^{\prime})-D_{\alpha}(Y_{0}\|Y_{0}^{\prime})\leq{\cal B}_{1}(\tau,\ell)+(E-1){\cal B}_{1}(0,\ell)+{\cal B}_{1}(0,T-E\ell)={\cal B}_{2}.$
@@
$D_{\alpha}(Y_{T}\|Y_{T}^{\prime})$ $\leq D_{\alpha}(Y_{E\ell}\|Y_{E\ell}^{\prime})+2\alpha[T-E\ell]\left(\frac{\zeta}{\sigma}\right)^{2}\leq D_{\alpha}(Y_{0}\|Y_{0}^{\prime})+2\alpha T\left(\frac{\zeta}{\sigma}\right)^{2}.$
Some remarks about Proposition A.5 are in order. First, part (a) shows that if $\phi_{t}$ and $\phi_{t}^{\prime}$ only differ at $t=1$, then $D_{\alpha}(Y_{T}\|Y_{T}^{\prime})$ is finite for any $T$. Second, part (a) also shows that if $\phi_{t}$ and $\phi_{t}^{\prime}$ differ cyclically for a cycle length of $\ell$, then the divergence between $Y_{T}$ and $Y_{T}^{\prime}$ grows linearly with the number of cycles $E$. Third, part (b) gives a bound that is independent of $L$. Finally, both of the bounds in parts (a) and (b) can be viewed as Rényi divergences between the Gaussian random variables ${\cal N}(0,\sigma^{2}I)$ and ${\cal N}(\mu,\sigma^{2}I)$ for different values of $\mu$.
In Appendix B, we give a detailed discussion of how the residuals $a$ from Lemma A.2 are chosen to prove Proposition A.5(a). In particular, we prove that the chosen residuals yield the tightest RDP bound that can achieved by repeatedly applying (20).
A.2 SGD operator analysis
This subappendix derives some important properties about the DP-SGD update operator $\cal{A}_{\lambda}(\cdot)$ in (12) and also contains the proof of Proposition 2.1.
To start, we recall the following well-known bound from convex analysis. Its proof can be found, for example, in [7, Theorem 5.8(iv)].
Lemma A.6.
if and only if
Lemma A.7.
$\left\langle\nabla f_{i}(x)-\nabla f_{i}(y),x-y\right\rangle\geq-m\left[1+\frac{m}{2(M+m)}\right]\|x-y\|^{2}+\frac{1}{2(M+m)}\|\nabla f_{i}(x)-\nabla f_{i}(y)\|^{2}.$
@@ A.2 SGD operator analysis
Define $F=f_{i}+m\|\cdot\|^{2}/2$ and let $x,y\in\operatorname*{dom}h$ be fixed. Moreover, note that $F$ is convex and satisfies (28) with $L=M+m$. It then follows from Lemma A.6 with $L=M+m$ that
$\leq\left\langle\nabla F(x)-\nabla F(y),x-y\right\rangle$
$\geq-\left[m+\frac{m^{2}}{2(M+m)}\right]\|x-y\|^{2}+\frac{1}{2(M+m)}\|\nabla f_{i}(x)-\nabla f_{i}(y)\|^{2}.$
Lemma A.8.
$\Delta:=u-v,\quad\Delta^{p}:=\operatorname{prox}_{\psi}(x)-\operatorname{prox}_{\psi}(y).$
Then, the following statements hold:
  • (a)
    $\|\Delta^{p}\|^{2}\leq\left\langle\Delta,\Delta^{p}\right\rangle$;
  • (b)
    $\|\Delta^{p}-\Delta\|^{2}\leq\|\Delta\|^{2}-\|\Delta^{p}\|^{2}$.
Proof.
(a) See [7, Theorem 6.42(a)].
(b) Using part (a), we have that
$\|\Delta^{p}-\Delta\|^{2}$ $=\|\Delta^{p}\|^{2}+\|\Delta\|^{2}-2\left\langle\Delta,\Delta^{p}\right\rangle\leq\|\Delta\|^{2}-\|\Delta^{p}\|^{2}.$
We now develop some technical properties of ${\cal A}_{\lambda}(\cdot)$. The first result presents a bound involving the following quantities for $x,y\in\operatorname*{dom}h$ and $i\in[k]$.
$\begin{gathered}d:=x-y,\quad\Delta_{i}:=\nabla f_{i}(x)-\nabla f_{i}(y),\quad\Delta_{i}^{p}:=\operatorname{prox}_{q_{i}}(\nabla f_{i}(x))-\operatorname{prox}_{q_{i}}(\nabla f_{i}(y)).\end{gathered}$ (29)
Lemma A.9.
Let $x,y\in\operatorname*{dom}h$ and $i\in[k]$ be fixed, let $d$, $\Delta_{i}$, and $\Delta_{i}^{p}$ be as in (29) for some $\{f_{i}\}$. Moreover, let $L_{\lambda}$ be as in (13), and suppose $f_{i}$ satisfies assumption (A2). If $\Delta_{i}^{p}=\Delta_{i}$, then for any $\lambda\leq 1/(M+m)$ we have
$\|d-\lambda\Delta_{i}^{p}\|\leq L_{\lambda}\|d\|.$ (30)
On the other hand, if $\Delta_{i}^{p}\neq\Delta_{i}$, then for any $\lambda\leq 1/[2(M+m)]$ we have
$\|d-\lambda\Delta_{i}^{p}\|\leq\sqrt{2}L_{\lambda}\|d\|.$ (31)
Proof.
Before proceeding, we first establish a technical inequality. Using Lemma A.7, it holds that for any $\mu>0$,
$\mu\left\|\Delta_{i}\right\|^{2}-2\left\langle d,\Delta_{i}\right\rangle$ $\leq\mu\|\Delta_{i}\|^{2}+2m\left[1+\frac{m}{2(M+m)}\right]\|d\|^{2}-\frac{\|\Delta_{i}\|^{2}}{M+m}$
$=2m\left[1+\frac{m}{2(M+m)}\right]\|d\|^{2}+\left({\mu-\frac{1}{M+m}}\right)\|\Delta_{i}\|^{2}.$ (32)
We now prove (30). Supposing that $\Delta_{i}^{p}=\Delta_{i}$, we have
$\|d-\lambda\Delta_{i}^{p}\|^{2}=\|d-\lambda\Delta_{i}\|^{2}=\|d\|^{2}+\lambda\left[\lambda\|\Delta_{i}\|^{2}-2\left\langle d,\Delta_{i}\right\rangle\right].$
Using (32) with $\mu=\lambda$, the above identity, and the definition of $L_{\lambda}$, it holds that for any $\lambda\leq 1/(M+m)$, we have
$\|d-\lambda\Delta_{i}^{p}\|^{2}$ $\leq\left(1+2\lambda m\left[1+\frac{m}{2(M+m)}\right]\right)\|d\|^{2}+\lambda\left({\lambda-\frac{1}{M+m}}\right)\|\Delta_{i}\|^{2}$
$=L_{\lambda}^{2}\|d\|^{2}+\lambda\left({\lambda-\frac{1}{M+m}}\right)\|\Delta_{i}\|^{2}\leq L_{\lambda}^{2}\|d\|^{2}.$
We now prove (31). Using Lemma A.8(b) with $(\Delta,\Delta^{p})=(\Delta_{i},\Delta_{i}^{p})$ and the inequality $\|a+b\|^{2}\leq 2\|a\|^{2}+2\|b\|^{2}$ for $a,b\in\mathbb{R}^{n}$, it holds that
$\left\|d-\lambda\Delta_{i}^{p}\right\|^{2}$ $=\left\|d-\lambda(\Delta_{i}+\Delta_{i}-\Delta_{i}^{p})\right\|^{2}\leq 2\|d-\lambda\Delta_{i}\|^{2}+2\lambda^{2}\|\Delta_{i}-\Delta_{i}^{p}\|^{2}$
[…]
$\leq 2\|d-2\lambda\Delta_{i}\|^{2}+2\lambda^{2}\|\Delta_{i}\|^{2}$
$=2\left(\|d\|^{2}+\lambda\left[2\lambda\left\|\Delta_{i}\right\|^{2}-2\left\langle d,\Delta_{i}\right\rangle\right]\right).$
Using (32) with $\mu=2\lambda$, the above inequality, and the definition of $L_{\lambda}$, it holds that for any $\lambda\leq 1/[2(M+m)]$, we have
$\|d-\lambda\Delta_{i}^{p}\|^{2}$ $\leq 2\left(1+2\lambda m\left[1+\frac{m}{2(M+m)}\right]\right)\|d\|^{2}+2\left({2\lambda-\frac{1}{M+m}}\right)\|\Delta_{i}\|^{2}$
$=2L_{\lambda}^{2}\|d\|^{2}+2\left({2\lambda-\frac{1}{M+m}}\right)\|\Delta_{i}\|^{2}\leq 2L_{\lambda}^{2}\|d\|^{2}.$
We are now ready to give the proof of Proposition 2.1.
A.2.1 Proof of Proposition 2.1
Proof.
$\leq\|x-y\|+2\lambda C.$
(b) Let $x,y\in\operatorname*{dom}h$ be arbitrary, and denote $\xi(\cdot):={\cal A}_{\lambda}(\cdot,\{f_{i}\},\{q_{i}\})$. Moreover, let $d$, $\Delta_{i}$, and $\Delta_{i}^{p}$ be as in (29). Using the fact that $\|\sum_{i=1}^{|B|}v_{i}\|^{2}\leq|B|\sum_{i=1}^{|B|}\|v_{i}\|^{2}$ for any $\{v_{i}\}\subseteq\mathbb{R}^{n}$, we have
$=\frac{1}{|B|^{2}}\left\|\sum_{i\in B}(d-\lambda\Delta_{i}^{p})\right\|^{2}\leq\frac{1}{|B|}\sum_{i\in B}\left\|d-\lambda\Delta_{i}^{p}\right\|^{2}.$ (33)
Using (33) and (31) in Lemma A.9, we conclude that
$\|\xi(x)-\xi(y)\|^{2}$ $\leq\frac{1}{|B|}\sum_{i\in B}\left\|d-\lambda\Delta_{i}^{p}\right\|^{2}\leq 2L_{\lambda}^{2}\|d\|^{2}=2L_{\lambda}\|x-y\|^{2}.$
(c) Let $\xi(\cdot)$, $d$, $\Delta_{i}$, and $\Delta_{i}^{p}$ be as in part (b). Following the same argument as in part (b), we obtain (33). Using (33) and (30) in Lemma A.9, we conclude that
$\|\xi(x)-\xi(y)\|^{2}$ $\leq\frac{1}{|B|}\sum_{i\in B}\left\|d-\lambda\Delta_{i}^{p}\right\|^{2}\leq L_{\lambda}^{2}\|d\|^{2}=L_{\lambda}\|x-y\|^{2}.$
A.3 RDP bounds
This subappendix derives the RDP bounds in Theorems 2.2 and 2.3.
The first result shows how the updates in Algorithm 1 are randomized proximal updates applied to the operator $A_{\lambda}(\cdot)$ in (12) with $q_{i}(\cdot)={\rm Clip}_{C}(\cdot)$.
Lemma A.10.
Let $\{X_{t}\}$, $\{X_{t}^{\prime}\}$, $\lambda$, $b$, $\sigma$, $C$, and $T$ be as in Theorem 2.2. Moreover, denote
$\phi_{t}(x):={\cal A}_{\lambda}(x,\{f_{i}\},\{{\rm Clip}_{C}\}),\quad\phi_{t}^{\prime}(x):={\cal A}_{\lambda}(x,\{f_{i}^{\prime}\},\{{\rm Clip}_{C}\}),\quad\forall x\in\operatorname*{dom}h,$
where ${\rm Clip}_{C}(\cdot)$ and ${\cal A}_{\lambda}(\cdot)$ are as in (2) and (12), respectively. Then, it holds that
Proof.
This follows immediately from the definition of $\phi_{t}$, the update rules in Algorithm 1, and the fact that ${\rm Clip}_{C}(\cdot)$ is the proximal operator of the (convex) indicator function of the convex set $\{x:\|x\|\leq C\}$. ∎
We now present some important norm bounds.
Lemma A.11.
Let $\{X_{t}\}$, $\{X_{t}^{\prime}\}$, $\{\phi_{t}\}$, and $\{\phi_{t}^{\prime}\}$ be as in Theorem 2.2 and denote $\ell=k/b$ and $t^{*}:=\inf_{t\geq 1}\left\{t:i^{*}\in B_{t}\right\}$. Then, it holds that
Proof.
$\leq\frac{\lambda}{b}\left[\|{\rm Clip}_{C}(\nabla f_{i^{*}}(x))\|+\|{\rm Clip}_{C}(\nabla f_{i^{*}}^{\prime}(x))\|\right]=\frac{2\lambda C}{b}.$
We now give the proofs of the main RDP bounds.
A.3.1 Proof of Theorem 2.2
Proof.
Using Proposition A.5(a) with
and $E=T=1$, we have that
$D_{\alpha}(X_{T}\|X_{T}^{\prime})$ $\leq D_{\alpha}^{(d_{h})}(X_{T-1}\|X_{T-1}^{\prime})+\frac{\alpha}{2\sigma^{2}}\left(Ld_{h}+\frac{2\lambda C}{b}\right)^{2}\theta_{L}(1)=\frac{\alpha}{2\sigma^{2}}\left(Ld_{h}+\frac{2\lambda C}{b}\right)^{2}.$
A.3.2 Proof of Theorem 2.3
Proof.
Suppose $f_{i^{*}}^{\prime}\neq f_{i^{*}}$ and $i^{*}\in B_{t^{*}}$ for indices $i^{*}$ and $t^{*}$. We first prove (17). In view of Proposition 2.1(b), it is clear that the DP-SGD update is $\sqrt{2}L_{\lambda}$-Lipschitz continuous. Hence, using Lemma A.11(b) and Proposition A.5(a) with
@@ A.3 RDP bounds
$=2\alpha\left(\frac{\lambda C}{b\sigma}\right)^{2}\left[E\theta_{L_{\lambda}}(\ell)+\theta_{L_{\lambda}}(T-t^{*}-1-E\ell)\right]\leq 2\alpha\left(\frac{\lambda C}{b\sigma}\right)^{2}\left[1+E\theta_{L_{\lambda}}(\ell)\right],$
where the last inequality follows from the fact that $\theta_{L_{\lambda}}(s)$ is nonincreasing for $s\geq 1$ and $\theta_{L_{\lambda}}(1)=1$.
We now prove (19). In view of (18) and Proposition 2.1(c), it is clear that the DP-SGD update, in the absence of gradient clipping, is $L_{\lambda}$-Lipschitz continuous. Consequently, the desired bound follows from the same arguments as in the proof of (19) above, but with $L=L_{\lambda}$ instead of $L=\sqrt{2}L_{\lambda}$. ∎